OpenStudy (anonymous):

Is the system consistent or inconsistent? y = -1/2x - 1 and y = 3x + 3 Consistent Inconsistent Cannot be determined I think it is inconsistent but im not to sure will someone help?

OpenStudy (anonymous):

Substitute the first y = ... in 2 nd equation. Then you get a quadratic in x. If it has real roots, then the system is consistent. Otherwise not.

OpenStudy (anonymous):

I agree with your answer, it is inconsistent because if, for instance, you solve for x you get x=8/7 and then the y value for the first equation is different to that of the second. Hence it is inconsistent

OpenStudy (anonymous):

thank you I always make simple mistakes I just wanted to be sure

OpenStudy (anonymous):

The system, 6x - 4y = 5 and 9x - 6y = 5, is consistent. ture/false

OpenStudy (anonymous):

I think it is true

OpenStudy (anonymous):

I think the first system is consistent, since they have a real solution.

OpenStudy (anonymous):

@amistre64

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

@jode0001 it is consistent?

OpenStudy (jdoe0001):

if it has a solution, yes

OpenStudy (anonymous):

okay explain please

OpenStudy (jdoe0001):

as koushik_ksv pointed out, is inconsistent when it has no solutions, or "dependent" when it has infiniite solutions

OpenStudy (anonymous):

o okay, I see